Circular No. 6063

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS J. Reynolds and D. Jauncey, Australia Telescope National Facility,

on behalf of the Southern Hemisphere Very Long Baseline Interferometry (SHEVE) team, report the results of 2.3-GHz VLBI and Parkes-Tidbinbilla Interferometer (PTI) observations of GRO J1655–40: “Over the 7-day period Aug. 17–24, the PTI and VLBI observations show the source to consist of a linear, highly-collimated structure dominated by two well-separated components. These components are moving apart at the rate of 0′′.095/day. At the 3.5-kpc minimum distance implied by AT Compact Array H I absorption observations, this corresponds to an apparent separation velocity of 1.9c and an overall size of 11.1 light-days on Aug. 21.3 UT. The observed angular size and expansion rate point to an origin on Aug. 15.5, which corresponds closely with the Aug. 15.2–15.3 early maximum seen at 843 MHz (see IAUC 6062 and 6055).”

R. J. Ivison, Royal Observatory, Edinburgh; and E. I. Robson, Joint Astronomy Centre, report: “The James Clerk Maxwell Telescope detected the x-ray transient GRO J1655–40 on Aug. 22.3 UT at the position reported on IAUC 6055. The 1.1- and 1.3-mm flux densities were 0.13 ± 0.02 and 0.13 ± 0.03 Jy, consistent with an extrapolation of the power law observed by the Very Large Array on Aug. 19.1 (IAUC 6060).”

H. Inoue, F. Nagase, M. Ishida, T. Sonobe, and Y. Ueda communicate, on behalf of the ASCA Team: “This object was observed with ASCA during Aug. 23.47–23.93 UT. The location of the x-ray source is α = 16^{h}53^{m}57^{s}.0, δ = −39^{o}50′55′′ (equinox 2000.0), close to the positions of the optical nova (IAUC 6050) and radio emission (IAUC 6055), within the systematic uncertainty of the ASCA image (error circle radius 1′). The flux in the range 1–10 keV changed by a factor of four, starting at 80 mCrab and increasing from 2 to 8 × 10^{−9} erg cm^{−2} s^{−1} during the observation. No apparent pulsation was found in the period range 1–1000 s. The spectrum in the range below 7 keV of the low-intensity state is represented by a power-law spectrum with photon index of ∼ 2.1 and interstellar absorption column density of 5 × 10^{21} cm^{−2}. An iron absorption edge is visible at 7.1 keV. An iron emission line around 6–7 keV and fine structures in the range 1–4 keV are possibly seen in the spectrum. The spectrum in the high-intensity state is similar but a little harder than that in the low-intensity state.”

1994 August 29 Daniel W. E. Green